C. Calculate the grams of lead (II) iodide that can be reproduced from 75.00 grams of potassium iodide


1
Expert's answer
2021-05-10T01:21:31-0400

75.00 g of KI = 75.00 / 166 = 0.452 moles


0.45 moles of KI produces 0.45 /2 = 0.226 mol PbI2


Amount of PbI2 produced = 0.226 x 461


= 104.1 grams